OpenStudy (anonymous):

helpppppppppppppp 25 5/8-17 15/16

OpenStudy (whpalmer4):

You need to get a common denominator to be able to subtract the fractions. An easy way to do that is to multiply both top and bottom (numerator and denominator) of each fraction by the denominator of the other fraction. Or, in this case, you might notice that one denominator is a multiple of the other (16 = 2*8). 25 + 5/8 - (17 + 15/16) = ? let's multiply 5/8 by 2/2 to get its value in 16ths: 25 + (2/2)*(5/8) = 25 + 10/16 now we can do our subtraction: (25 + 10/16) - (17 + 15/16) = (25-17) + (10/16-15/16) = 8 - 5/16 or 7 + 16/16 - 5/16 = 7 11/16. If we had something like 2/3 - 1/4, we would multiply 2/3 by (4/4) and 1/4 by (3/3) to get (2/3)(4/4) - (1/4)(3/3) = 8/12-3/12 = 5/12.